what is the new volume of a 5.0 L sample of nitrogen gas that is heated from 95 degrees celceious to 170 degrees celceious
1
Expert's answer
2012-09-13T12:10:40-0400
P1V1/T1=P2V2/T2
P = const
v1/t1=v2/t2
v2=v1*t2/t1= 5*
(273+170)/(273+95)=6.02 L